A mother is 30 years old. Her daughter is 4 years old. In how many years will the age of the mother be triple the age of her daughter?
I've tried doing it by trial and error but I am looking for the equation as to how I can find this.
Thank you!

Answer by Alan3354(69443)

A mother is 30 years old. Her daughter is 4 years old. In how many years will the age of the mother be triple the age of her daughter?
M = D + 26
---------
In x years,
M + x = 3*(D + x)
Sub for M
D+26 + x = 3D + 3x
D = 4
30 + x = 12 + 3x
2x = 18
x = 9
---> in 9 years